OpenStudy (blackops2luvr):

The leg of a right triangle is 5 units and the hypotenuse is 8 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (blackops2luvr):

@iGreen

OpenStudy (igreen):

Use the pythagorean theorem: \(a^2 + b^2 = c^2\) \(a^2 + 5^2 = 8^2\) \(a^2 + 25 = 64\) Subtract 25 to both sides: \(a^2 = 39\) \(a = \sqrt{39}\) Can you square that? @blackops2luvr

OpenStudy (igreen):

@blackops2luvr

OpenStudy (blackops2luvr):

Answer is sqrt 39 correct?

OpenStudy (igreen):

Yep^^

OpenStudy (blackops2luvr):

okay.

OpenStudy (blackops2luvr):

I poasted it on Chatstep.

OpenStudy (blackops2luvr):

okay. here.

OpenStudy (igreen):

Ehem, post a new question :P

OpenStudy (blackops2luvr):

lol okay.